### v2

# iCost: A Novel Instance Complexity Based Cost-Sensitive Learning Framework

Abstract
Class imbalance in data presents significant challenges for classification tasks. It is fairly common and requires careful handling to obtain desirable performance. Traditional classification algorithms become biased toward the majority class. One way to alleviate the scenario is to make the classifiers cost-sensitive. This is achieved by assigning a higher misclassification cost to minority-class instances. One issue with this implementation is that all the minority-class instances are treated equally, and assigned with the same penalty value. However, the learning difficulties of all the instances are not the same. Instances that are located in the overlapping region or near the decision boundary are harder to classify, whereas those further away are easier. Without taking into consideration the instance complexity and naively weighting all the minority-class samples uniformly, results in an unwarranted bias and consequently, a higher number of misclassifications of the majority-class instances. This is undesirable and to overcome the situation, we propose a novel instance complexity-based cost-sensitive approach (termed ’iCost’) in this study. We first categorize all the minority-class instances based on their difficulty level and then the instances are penalized accordingly. This ensures a more equitable instance weighting and prevents excessive penalization. The performance of the proposed approach is tested on 65 binary and 10 multiclass imbalanced datasets against the traditional cost-sensitive learning frameworks. A significant improvement in performance has been observed, demonstrating the effectiveness of the proposed strategy.
keywords:
Cost-Sensitive Learning , Imbalanced Classification , Multiclass classification , Data difficulty , Scikit-Learn , Weighted classifier
{highlights}
Cost-sensitive learning is a popular technique used in the imbalanced domain.
Higher misclassification costs are naively applied to all minority-class instances.
In the proposed approach, instances are penalized according to their complexity.
This ensures a plausible weighting mechanism, resulting in fewer misclassifications.
The iCost algorithm, written in Python, is fully compatible with the scikit-learn library.
1 Introduction
When the class distribution in the dataset is uneven, with one class significantly outnumbering the other, it is referred to as imbalanced data [1]. Class imbalance in data is frequently encountered in different applications such as medical diagnosis, fraud detection, spam detection, etc. [2]. This can pose significant challenges for standard machine learning (ML) algorithms, which typically assume that the classes are balanced. Consequently, ML classifiers produce biased performance towards the majority class. Imbalanced learning is a critical area in ML that requires specialized techniques to ensure that models are effective and fair, especially in applications where the cost of misclassifying minority instances can be dire [3]. As such, the imbalanced domain has caught a lot of attention from researchers and different approaches have been proposed to address the issue [4]. The techniques can be broadly classified into two categories: data-level approach and algorithmic-level approach.
In the data-level approach, the original class distribution in the data is modified by adding new synthetic minority-class instances or eliminating samples from the majority class [5]. The goal is to balance the class distribution in the data. Recent research suggests it is even more important to reduce the class overlapping in the process to obtain better performance [6]. On the other hand, in the algorithmic-level approach, the original classification algorithm is modified to adapt to the imbalanced domain scenario. This is achieved by changing the cost function to handle the class imbalance directly [7]. Higher misclassification costs are assigned to the minority class instances to make the algorithm more sensitive to those errors. During training, the model learns by trying to reduce the overall misclassification cost. Assigning higher weight to the minority-class misclassifications shifts the bias from the majority class. This way, the algorithm is made cost-sensitive (CS). This approach is classifier-dependent as different algorithms use different learning procedures. Both of these approaches perform almost equally well and are extensively used in different real-world applications [8, 9, 10, 11].
This study is focused on cost-sensitive learning (CSL). Here, a specific penalty is added to the misclassifications of the minority-class instances. Standard classifiers use a 0-1 loss function for calculating cost. This indicates a value of 0 for correct classifications and a value of 1 for incorrect classifications. This type of error-driven (ED) classifier assumes an even class distribution in the data. However, when the data is skewed, this approach does not fare well and fails to achieve high sensitivity (accuracy of the minority class prediction). This happens because misclassifications of both majority and minority class instances are given equal weight. In many applications, correctly classifying the minority-class instances, which usually represent the positive cases, is more important [9]. Therefore, the idea of the cost-driven classifier is introduced, where asymmetric misclassification cost is utilized. Assigning a higher misclassification cost to the minority-class instances compared to the majority-class forces the algorithm to put more priority on learning those instances correctly, reversing the bias. This approach works quite well when the data is imbalanced and has been incorporated into the implementation of different classification algorithms in scikit-learn and similar libraries such as xgboost [12, 13].
Table 1: Cost Matrix
Predicted True Predicted False
Actual True 0 Cr Minority Class
Actual False Cp 0 Majority Class
[…]
The effect of modifying the weights of the instances on the decision boundary for the Support Vector Machine (SVM) classifier is illustrated in Fig 1. Assigning a higher weight to some instances forces the classifier to put more emphasis on getting those points right, deforming the original decision boundary. This often leads to misclassifications of some of the nearby unweighted samples, as can be observed from the figure. Therefore, instances should be weighted very carefully to avoid unusual predicaments.
Figure 1: Effect of modifying weights on the decision boundary. Here, the size of points is proportional to its weight.
In the case of the traditional CS approach, the assigned cost value is applied to all the minority class instances indiscriminately and this raises a major concern [14]. All minority class instances do not pose the same level of difficulty. Samples that are closer to the decision boundary have a higher chance of getting misclassified than those that are far away. The more difficult-to-learn samples should be penalized more heavily than the others. This instance-level difficulty characteristic has not been considered in previous literature and our study addresses this issue. For instance, let’s assume a dataset has an IR of 100. Then all the minority-class instances will be penalized 100 times more strongly than any majority-class instances. This will create some unusual deformation of the decision boundary and consequently, cause a higher number of misclassifications of the majority class instances (lower specificity score) during testing. Assigning unnecessarily high penalties to the minority-class samples that are located far away from the decision boundary only biases the predictions toward the minority-class. Consequently, some nearby majority-class instances get wrongly classified. Overall, the algorithm struggles to perform well on the test data. To mitigate the situation, an instance-difficulty-based cost should be applied. We propose such an algorithm termed ”iCost: Instance complexity-based Cost-sensitive learning” in this study.
Now, how to faithfully define the complexity of the instances remains an issue. Class overlapping between different class instances is usually identified as the most crucial factor that affects the classifier’s performance [15]. The presence of noisy samples in the data is another major concern. Small disjunct in the dataset also makes the learning task difficult for the algorithms [16]. These issues have been illustrated in Fig. 2 and Fig. 3. As we can see from figure 3, there are two minority-class instances completely surrounded by majority-class samples. These two samples are likely outliers as they are far away from other minority-class instances. However, in the traditional CS approach, these two noisy samples will also be penalized. This will obviously cause misclassifications of some of the nearby majority-class instances. Moreover, the model will fail to generalize well in unseen test data. As for figure 2, the instances in the overlapping region are more difficult to classify as they are surrounded by some opposite-class instances. These particular samples should be penalized more stringently than instances in the non-overlapping region. We consider all these parameters while defining instance complexity for our proposed algorithm.
Figure 2: Class overlapping between opposite class instances
Figure 3: Presence of noisy instance in the data (red circle marks the noisy minority-class instances). It also illustrates small disjunct in the data.
Researchers have proposed various ways of identifying overlapping regions. A taxonomy of different class overlap measures was presented by Santos et al. [17]. The authors grouped the approaches into four different categories: feature overlap, instance overlap, structural overlap, and multiresolution overlap. Each category of approach undertakes a different method for quantifying class overlap. For instance, maximum Fisher’s discriminant ratio (F1) is one of the most widely used feature-based measures to compute class overlap [18]. Clustering methods can also be utilized to determine the degree of overlap [19]. In this study, we used two different methodologies to define instance complexity. One is based on a neighborhood search algorithm. The other one uses concepts from graph theory by analyzing information on the internal structure of classes. These approaches allow us to take into consideration the class overlapping problem as well as the presence of noisy instance scenarios. In the future, we plan to incorporate a few other methods such as hypersphere coverage [20], and local set [21] to define instance complexity.
In the first approach, a neighborhood search algorithm is employed to grade the samples into different levels. This is done based on the number of neighboring samples from the opposite class. A higher misclassification cost is assigned to the samples that are surrounded by more instances of the opposite class. This method also allows us to identify noisy instances in the data. This way, minority-class samples in the overlapped regions are prioritized significantly over the majority-class samples, leading to correct identifications of those hard-to-learn instances. On the other hand, samples that are in the safe zones surrounded by samples from the same class are marginally weighted, reducing their effect. In the second approach, a Minimum Spanning Tree (MST) is constructed over the data to determine structural overlap. The minority-class instances that are connected to the opposite class by an edge in the MST in considered boundary samples and therefore highly penalized compared to others. This way, asymmetric cost is applied to different minority-class instances based on their complexity. This ensures a more appropriate distribution of weights among the minority-class instances. This not only helps in reducing the majority-class misclassifications but also improves the performance of the minority-class. Thus, our proposed algorithm is able to provide improved prediction performance on a wide range of imbalanced datasets.
The appropriate value for the penalty can be determined using grid search, evolutionary techniques, or similar approaches. Although these techniques are likely to enhance predictive performance, excessive optimization may overfit the data. The algorithm code is written in Python. It is implemented by inheriting classifiers from the sklearn library, ensuring full compatibility with sklearn’s CS classifiers. Different ways of categorization of the minority-class instances are possible and some of them have been included in our implementation. The performance of the proposed approach is evaluated on 65 binary imbalanced datasets. The algorithm is further extended for multiclass scenarios and tested on 10 multiclass imbalanced datasets. Significant performance improvements have been observed as compared to traditional CS approaches (employed in sklearn implementation) as well as some popular data resampling techniques. The results obtained affirm the advantage of the proposed approach.
The rest of the article is organized as follows. Related works have been discussed in Section II. The proposed methodology is presented in Section III. In Section IV, we laid out the experimental setup. The performance measures are reported in Section V. The manuscript concludes with Section VI, with a summary of the work. For reproducibility and the benefit of the community, all codes and results are made available in the following GitHub repository: https://github.com/newaz-aa/iCost.
2 Related Works
A lot of different techniques have been proposed over the years to deal with imbalanced data [22, 23]. However, very few of them address data-intrinsic characteristics [24]. Several data difficulty factors have been identified that are primarily responsible for making imbalanced classification so difficult [16]. This includes class overlapping, small disjuncts, noisy samples, and rarity of samples among others. It has been suggested that the overlapping between the classes, not the imbalance, is primarily responsible for complicating the learning task [25]. Recent literature emphasizes the importance of considering data difficulty factors to develop more robust methods for learning from imbalanced datasets [16]. Stefanowski proposed several approaches to tackle these data difficulty factors [26]. However, the study only considers data resampling techniques. In this study, we are primarily focused on CS approaches.
Different classifiers such as SVM, Artificial Neural Network (ANN), and Decision Tree (DT) are adapted to the CS framework [27, 28]. Their implementation is available in the popular sklearn library. In different medical datasets, class imbalance is prevalent and CS approaches are found to be quite successful in handling such scenarios [29]. Mohosheu et al. performed a detailed efficacy analysis of the performance of CS approaches as compared to data resampling techniques in their study [30]. The authors reported that the traditional CS approaches outperform undersampling and ensemble methods but cannot surpass popular oversampling techniques such as SMOTE or ADASYN [31]. Hybridization between sampling and CS approaches is also possible [32, 33]. The authors first resampled the data by generating synthetic samples using SMOTE and then used a weighted XGBoost classifier for training. The authors suggested that combining the techniques allows for lower misclassification costs to be applied while also reducing the number of synthetic samples to be generated. Consequently, overfitting is reduced and better prediction performance can be obtained.
Natarajan et al. explored two different ways to improve binary classification performance using CSL under noisy labels, focusing on real-world applications. The authors also reported that weighted LR and SVM classifiers are somewhat noise-tolerant. Gan et al. proposed a sample distribution probability-based CS framework in their article [34]. Roychoudhuri et al. adapted the CS algorithm for time-series classification [35]. Zhou et al. extended the CS framework for multiclass imbalanced scenarios [36]. Other variations of CS approaches include MetaCost [37], a meta-learning algorithm that converts any given classifier into a CS classifier. The idea of example-dependent cost has also been proposed in previous literature [38, 39]. For instance, in credit scoring, a borrower’s credit risk is determined based on different factors which include their credit history and financial behaviors. These factors should be taken into consideration while weighting instances for predictive modeling [40]. However, such approaches are application-specific and do not generalize well for other datasets. A detailed review of different CS methods has been presented in this recent article [41]. None of these CS approaches considers instance-difficulty-based characteristics. This has been pointed out in some recent literature [6, 14, 16] and we address this issue in our study.
3 Methodology
3.1 Cost Sensitive Classifier
In classification algorithms, a cost function is utilized for the learning process. The goal is to minimize the total cost associated with the predictions. Optimization algorithms such as stochastic gradient descent are utilized to reach the optimal point. This way, the model parameters are determined and the classifier learns to predict as accurately as possible. To give an example, for the LR classifier, log loss is calculated in the following way -
Log $\text{Log Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(h_{\theta}(x_{i}))+(1-y_{i})\log(1-h_{\theta}(x_{i}))\right]$
Here,
  • 1.
    N is the total number of samples
  • 2.
    $y_{i}$ is the true label of the ith sample (0 or 1)
  • 3.
    $h_{\theta}(x_{i})$ is the predicted probability that the ith sample belongs to class 1, as calculated by the classifier.
  • 4.
    $\theta$ is the model parameters.
The first term in the sum represents the cost for the positive class and the second term represents the cost for the negative class. Now, for the correct predictions, the cost becomes 0. As for the incorrect predictions, the cost is calculated for all misclassifications. Here, no distinction is made between the positive and negative instances. All the misclassifications are weighted equally. To make the classifier cost-sensitive, the log-loss formula can be modified by introducing different weights.
CS Log $\text{CS Log Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[C_{1}\cdot y_{i}\log(h_{\theta}(x_{i}))+C_{0}\cdot(1-y_{i})\log(1-h_{\theta}(x_{i}))\right]$
Here,
  • 1.
    $C_{1}$ is the cost of misclassifying a positive example.
  • 2.
    $C_{0}$ is the cost of misclassifying a negative example.
This formula allows adjusting the penalty for each class by tuning $C_{1}$ and $C_{0}$. This is the traditional CS approach where different weights are assigned to majority and minority class misclassifications. Usually, the $C_{0}$ value is kept fixed at 1 while the $C_{1}$ is assigned a different weight (typically, the IR of the dataset). This allows the algorithm to put more emphasis on correctly predicting the minority-class instances. Thus, modifying the learning process. While this approach certainly improves the model’s ability to correctly identify the minority-class instances, this usually comes at the cost of many misclassifications of the majority-class instances as they are assigned a lower weight. The trade-off becomes a defining factor here that how many misclassifications of the majority-class instances can be justified in exchange for a few correct predictions of minority-class instances. Although in certain applications such as medical diagnosis or fault detection, correctly identifying the positive cases is crucial, however, too many false positives reduce the reliability of the prediction framework.
In our proposed methodology, we try to address this critical but often neglected issue through an intuitive hypothesis. In the traditional CS framework, an underlying issue is ignored. All the instances in the dataset do not have the same difficulty of learning. Depending on their location in the feature space, some are easier to classify while others can be difficult. Instances that are closely located i.e., share similar feature attributes can be difficult to distinguish from one another. Depending on such characteristics, we propose to divide the instances into multiple categories. Then each category of samples is weighted differently in accordance with their complexity level. Mathematically, it can be represented as follows -
Modified CS Log $\text{Modified CS Log Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[C_{j(i)}\cdot y_{i}\log(h_{\theta}(x_{i}))+(1-y_{i})\log(1-h_{\theta}(x_{i}))\right]$
Here, $C_{j}(i)$ is the cost for the category j that instance i belongs to. The concept of categorization of the instances is discussed in the following section. One thing to consider here, only the minority-class instances (positive cases) are categorized and penalized with different weights. The majority-class instances are assigned equal weights, in line with the traditional cost-sensitive approach. While majority-class instances can also be categorized in a similar fashion, that would increase the complexity of the algorithm and is not explored in this article.
This concept is equally applicable to other classification algorithms such as SVM or DTs. For the SVM classifier, the cost (objective) function aims to find a hyperplane that maximizes the margin between classes while penalizing misclassified points. For a linear SVM, the objective function can be expressed as follows:
$\text{Minimize}\quad\frac{1}{2}||\mathbf{w}||^{2}+C\sum_{i=1}^{N}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))$
Here,
  • 1.
    w is the weight vector defining the hyperplane.
  • 2.
    $||w||^{2}$ is the regularization term that controls the margin width.
  • 3.
    C is the penalty parameter that controls the trade-off between maximizing the margin and minimizing classification errors.
  • 4.
    $y_{i}$ is the true label for instance i.
  • 5.
    $x_{i}$ is the feature vector for instance i.
  • 6.
    b is the bias term.
  • 7.
    $max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b)$ is the hinge loss, penalizing instances on the wrong side of the margin.
The function can be modified to better handle imbalanced data by penalizing errors on minority-class instances more heavily. The objective function for the CS-SVM classifier can be expressed as follows:
$\text{Minimize}\quad\frac{1}{2}||\mathbf{w}||^{2}+C_{+}\sum_{i\in\mathcal{P}}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))+C_{-}\sum_{i\in\mathcal{N}}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))$
Here,
  • 1.
    $C_{+}$ is the penalty parameter for the positive class (often the minority class).
  • 2.
    $C_{-}$ is the penalty parameter for the negative class (often the majority class).
  • 3.
    $\mathcal{P}$ and $\mathcal{N}$ represent the sets of positive and negative instances, respectively.
We can incorporate the proposed instance-based weighting formula into the equation in the following way -
$\text{Minimize}\quad\frac{1}{2}||\mathbf{w}||^{2}+C\sum_{i\in\mathcal{M}}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))+\sum_{j=1}^{k}C_{j}\sum_{i\in\mathcal{C}_{j}}\max(0,1-y_{i}(\mathbf{w}\cdot\mathbf{x}_{i}+b))$
Here,
  • 1.
    C is the penalty parameter for all majority-class instances, represented by set $\mathcal{M}$.
  • 2.
    $C_{j}$ is the penalty parameter specific to category j within the minority class.
  • 3.
    k is the number of categories in the minority class.
This way, the cost-sensitive learning framework of different classifiers can be adjusted to incorporate the instance-complexity-based weighting mechanism into the algorithm.
3.2 Instance Complexity
In this manuscript, we present two distinct approaches to defining instance complexity: one based on a neighborhood search algorithm and the other on a graph algorithm.
3.2.1 Neighborhood Search
When a minority-class instance is surrounded by majority-class samples, it naturally becomes difficult to classify them correctly. In the nearest neighbor rule, if there are more instances from the opposite class surrounding the given example, then it would be misclassified. We utilize this concept to identify hard-to-learn examples in the data. At first, we identify the K-nearest neighbors of each minority class sample. A k-value of 5 was utilized in this study. The nearest neighbors are computed using the Euclidean distance. Next, each minority-class instance is categorized as follows:
  • 1.
    Pure: Number of neighboring samples belonging to the majority class = 0
  • 2.
    Safe: Number of neighboring samples belonging to the majority class = 1 or 2
  • 3.
    Border: Number of neighboring samples belonging to the majority class $>2$
This has been illustrated in Fig. 4. Samples that are categorized as ’pure’ are completely surrounded by instances of the same class. These samples are easy to classify and are usually located far from the decision boundary. ’Safe’ samples have 1 or 2 neighboring opposite-class instances. These samples should be handled carefully due to the risk of misclassification. Border samples are surrounded by majority-class instances. These samples would be misclassified by the K-nearest neighbor classification rule.
This categorization formula is flexible and can be easily modified. We keep that option in the implementation of the algorithm.
Figure 4: Categorization of Minority-class instances
3.2.2 Minimum Spanning Tree (MST)
A spanning tree is a tree-like subgraph of a connected, undirected graph that contains all the vertices of the graph. A minimum spanning tree shares the characteristics of a spanning tree but with the additional condition of having the smallest total edge weights among all possible spanning trees. In this approach, the MST is first built on the dataset by connecting all the instances. Next, all the vertices (data points) that have an edge connecting them from 2 distinct classes are identified. The minority-class instances that are connected to the opposite class by an edge of the MST are penalized heavily compared to other minority-class samples.
This has been illustrated in Fig. 5. In this approach, the minority-class instances are divided into only two categories. While this offers less flexibility, it is easily tunable and provides better performance than the traditional CS approach. Here, unlike the neighborhood approach, the algorithm focuses on the immediate neighbor and penalizes accordingly. The Scipy library was utilized to calculate the MST.
Figure 5: Categorization of Minority-class instances based on MST. The marked points represent linked examples.
3.3 Proposed Algorithm
The proposed approach can be broadly divided into two segments: categorization of the minority-class instances and their weighting.
3.3.1 Minority-class instance categorization
In the first stage. the minority-class instances are categorized based on their complexity. Two criteria have been implemented so far, as described in the previous section.
In the neighborhood search approach, instances are categorized into three 3 distinct groups. For the ’pure’ group, a comparatively much smaller misclassification cost should be enough to correctly identify these samples. Assigning a higher weight might worsen the scenario by unusually deforming the decision boundary. This may result in a higher number of misclassifications of any nearby majority-class instances. For the ’safe’ group, assigning too small a weight may be insufficient, while too large a weight could reverse the situation. On the other hand, for the ’border’ group, a higher weight is necessary to prioritize these samples over the neighboring samples from the opposite class.
In the MST approach, instances are categorized into two distinct groups. The instances that are linked to the opposite class are strongly penalized. These instances are usually located along the border. The remaining instances are marginally penalized. The weight of majority class misclassification is usually kept as 1 for simplicity. However, for the linked majority-class instances, the misclassification cost can be reduced to minimize their effect.
3.3.2 Instance Weighting
In our implementation, different weights are assigned to different categories of samples based on their difficulty level. We employed a grid-search technique to determine the appropriate cost for different categories during experiments. The values varied from one dataset to another. Based on the experiments conducted on 66 datasets, we set a default value for each category for direct implementation. The default setting of the weights also provides quite an improvement in performance. Employing different search algorithms such as evolutionary techniques to find the more suitable weights for individual datasets can further optimize the prediction performance.
The default penalty value for border samples is fixed as the IR of the dataset. For the safe samples, half of that value is utilized. For the ’pure’ category, a cost factor of 0.2*IR is selected which is usually closer to the misclassification cost of majority class samples.
The default penalty value for linked instances is set to the IR of the dataset. As for the other samples, half of IR is used as the misclassification cost.
3.3.3 Additional Features
There are other ways of categorizing minority-class instances. For instance, we can include another category ’outlier’ for those instances that are completely surrounded by opposite classes. The associated penalty for these instances should be limited to avoid overfitting. This can also implemented in our proposed framework by assigning four different penalties. However, employing search algorithms to find a suitable set of four different values can be a bit more time-consuming. While this was not tested in the experiment, we hypothesize that it could prove highly effective in large datasets with many outliers.
To give the user this freedom, we provide a general categorization formula based on the number of majority-class samples surrounding a minority-class instance. We grade each minority-class sample from g0 to g5 based on 0 to 5 neighboring majority-class samples, respectively. Then, different weights can be assigned to each type of minority-class sample based on its grade. It also opens for some other considerations, such as the presence of outliers. The minority-class samples that are labeled as ’g5’ are completely surrounded by instances of the opposite class. One possible explanation for this could be that these are noisy instances within the majority class. Assigning higher weight to such instances may lead to the misclassification of several neighboring majority-class samples. In the proposed methodology, we have the option to ignore such instances ensuring better generalization.
3.4 Implementation
The proposed approach is implemented using the Python programming language. It is built on top of the scikit-learn library. It inherits from two parent classes: BaseEstimator and ClassifierMixin, to ensure compatibility with scikit-learn’s tools, allowing it to integrate into ML pipelines, perform hyperparameter tuning, and use standard methods like fit, predict, and score.
The architecture is presented below:
Algorithm: Instance complexity-based Cost-sensitive learning (iCost)
Inputs:
  • 1.
    data: Input dataset (Pandas DataFrame object)
  • 2.
    base_classifier: A classifier instance (e.g., LogisticRegression(), SVC(), RandomForestClassifier(), etc.). The framework inherits from sklearn’s implementations of those algorithms.
  • 3.
    algorithm: ’original’, ’neighborhood’ (default), or ’MST’. Here, ’original’ refers to the traditional implementation of the CS classifiers where all the minority-class instances are weighted equally.
  • 4.
    type: ’ins’ or ’gen’ (default = ’ins’).
    • (a)
      ’ins’ refers to the instance categorization criteria (3 groups) we proposed.
    • (b)
      ’gen’ provides the general categorization (g0 to g6) mentioned earlier.
    These are only applicable if the ’algorithm’ parameter is set to ’neighborhood’.
  • 5.
    n_neighbors: The number of neighbors to be considered for categorization of the minority-class instances under ’neighborhood’ criteria. The default value is set to 5 throughout the experiment.
  • 6.
    cost-factor: The misclassification costs to be assigned. It can be an integer, a list, or a dictionary. This input parameter is related to the ’algorithm’ and ’type’ parameters.
    • (a)
      For algorithm = ’original’, the cost-factor value must be an integer. The default is set to the IR of the dataset.
    • (b)
      For algorithm = ’neighborhood’ and type = ’ins’, there are three different weight parameters: cfb, cfs, and cfp. These refer to the costs associated with the border, safe, and pure samples, respectively. Evidently, $cfb>cfs>cfp$. The default values are set as IR, 0.5*IR, and 0.25*IR, respectively.
    • (c)
      For algorithm = ’MST’, there are two different weight parameters: cfl, and cfn. These refer to the costs associated with the linked and normal samples, respectively. The default values are set as IR and 0.5*IR, respectively.
Output: Instance-level weighted classifier object.
Constraints: The base classifier must support the ”sample_weight” parameter. This enables the classifiers to be made cost-sensitive. Certain classification algorithms (Naive Bayes, for instance) are not inherently designed for cost-sensitive classification. So, they cannot be used as a base classifier.
Procedure:
  • 1.
    If algorithm = ’original’, the algorithm assigns a weight equal to the cost factor to all the minority-class instances without any other consideration. This is the original CS implementation of the algorithms. If the cost factor value is 1, the algorithm will work as a standard ED classifier.
  • 2.
    If type = ’ins’, the algorithm categorizes the minority-class instances into three categories. For ’gen’, minority-class instances are categorized into k+1 categories.
  • 3.
    In the case of ’ins’, the user can provide an integer, an array, or a dictionary with three elements as the input values for the cost factor. If it is an integer, then the following penalty values are assigned.
    • (a)
      ’border’ samples: a penalty equal to the integer is assigned.
    • (b)
      ’safe’ samples: half of the integer value is assigned.
    • (c)
      ’pure’ samples: 0.25 times the integer value is assigned.
    If the input is an array, values are directly assigned to border, safe, and pure samples, in that order. In the case of a dictionary, key-value pairs can be used to directly state the cost values for each pair.
    • (a)
      list: cost-factor = [10, 4, 1.5]
    • (b)
      dictionary: cost-factor = {’cfb’: 10, ’cfs’: 4, ’cfp’: 1.5}
  • 4.
    The same thing goes for algorithm = ’MST’. Here, the minority-class instances are categorized into two distinct groups: linked and normal.
    If the input value for the cost factor is an integer, then the weights are assigned to the minority class samples in the following way:
    • (a)
      linked = cost factor
    • (b)
      normal = 0.5 * cost factor
    The user can also directly assign weights using an array or dictionary with two elements.
  • 5.
    For ’gen’, the user can assign weights using an array of k+1 elements. In the case of integer input or default scenario (weight=IR), the weight is equally divided between the samples from 1 to IR proportionally based on their grade. A dictionary can also be utilized with keys being grades g0 to g6.
Other Considerations: There are a few other issues that should be considered while applying the algorithm. They are discussed below.
  • 1.
    The appropriate values for the misclassification cost are dependent on the dataset. In the implementation of the sklearn library, a penalty equal to the IR of the dataset is employed. However, that value does not necessarily produce the optimal result. We performed experiments with different cost values on several imbalanced datasets and observed that the optimal value usually lies close to the IR of the dataset.
    In the proposed algorithm, the minority-class instances are divided into three groups which require three different cost values. Experiments were conducted with different cost settings on several imbalanced datasets and based on the observation, the default values are set for border, safe, and pure category samples. These default cost settings provide a decent improvement in performance over the traditional CS classifiers.
    However, the performance can be further optimized using different search algorithms such as grid search, random search, bayesian optimization, evolutionary algorithms, etc. [42]. The cost factor is the most sensitive parameter in our proposed framework and we suggest tuning it for optimal results.
  • 2.
    Any misclassifications of the majority class samples are assigned a weight of 1. Therefore, assigning a weight lower than 1 to any minority class instance can result in poor sensitivity in imbalanced classification tasks. Since minority-class samples are usually more important to classify correctly, a conditional statement is kept to ensure that the minimum weight assigned to any minority-class instance should not be lower than 1.
Example:
  • 1.
    iCost(base_classifier = LogisticRegression(), algorithm = ’neighborhood’, type = ’gen’, cost-factor = [2, 5, 5, 10, 10, 1])
  • 2.
    This will apply an instance complexity-based cost-sensitive LR classifier on the given data. Here,
    • (a)
      g0 graded minority-class samples (’pure’ category) are weighted by a factor of 2.
    • (b)
      The g1 and g2 graded minority-class samples (’safe’ category) are weighted by a factor of 5.
    • (c)
      The g3 and g4 graded samples are weighted by a factor of 10.
    • (d)
      The remaining samples (g6 category) which are entirely surrounded by 5 opposite-class instances are more likely to be outliers and penalizing such instances highly can cause overfitting on minority-class data. Therefore, those are weighted by a factor of 1.
  • 3.
    While such intuitive settings of the cost factor parameter are likely to improve the prediction performance, it is better to tune the parameter to find the most appropriate cost settings and obtain optimal results.
3.5 Extension - Multiclass Classification
The proposed algorithm can be applied for multiclass classification as well. Compared to binary classification scenarios, multiclass classification is much more complicated and requires more sophisticated strategies [43].
Classification through binarization i.e. decomposition strategies such as One-vs-One (OvO) or One-vs-Rest (OvR) are the most commonly used approach used in multiclass classification. In these approaches, the multiclass problem is disintegrated into multiple binary classification problems. These methods have their own drawbacks. For instance, in OvR, even when the classes are balanced, the binarization technique forces the data to become skewed, making the classification more complex [44]. However, these techniques are quite effective in handling data in multiple classes.
The sklearn library supports such decomposition techniques through classes such as OneVsRestClassifier and OneVsOneClassifier. These classifiers take an estimator object as the input. We also ensured the compatibility of our proposed framework with these approaches. To apply the proposed framework in multiclass scenarios, the iCost object just needs to be fed as the estimator to those classifiers. For the experiment, we utilized the OneVsRestClassifier with the 4 base classifiers as estimators to obtain the results.
4 Experiment
4.1 Datasets
The performance of the proposed algorithm has been evaluated on datasets with varying degrees of imbalance to ensure the generalizability of the proposed approach. The datasets are collected from KEEL and the UCI data repository [45]. All the datasets are publicly available. A handful of them have a few missing entries. The missing values are imputed using the MICE algorithm [46]. A total of 65 binary classification datasets and 10 multiclass classification datasets were utilized in this study. A summary of the datasets is provided in Table 2 and Table 3.
Table 2: Summary of the datasets for binary classification tasks
Dataset Name # Samples # Features Imbalance Ratio
glass1 213 10 1.8
wisconsin 683 10 1.86
pima 768 9 1.87
glass0 213 10 2.09
yeast1 1483 9 2.46
vehicle2 846 19 2.88
vehicle1 846 19 2.9
vehicle3 846 19 2.99
vehicle0 845 19 3.27
new-thyroid1 215 6 5.14
ecoli2 336 8 5.46
glass6 214 10 6.38
yeast3 1484 9 8.1
yeast 1484 9 8.1
ecoli3 336 8 8.6
page-blocks0 5472 11 8.79
ecoli-0-3-4_vs_5 200 8 9
yeast-2_vs_4 514 9 9.08
ecoli-0-6-7_vs_3-5 222 8 9.09
ecoli-0-2-3-4_vs_5 202 8 9.1
yeast-0-3-5-9_vs_7-8 506 9 9.12
glass-0-1-5_vs_2 172 10 9.12
yeast-0-2-5-7-9_vs_3-6-8 1004 9 9.14
yeast-0-2-5-6_vs_3-7-8-9 1004 9 9.14
ecoli-0-4-6_vs_5 203 7 9.15
ecoli-0-2-6-7_vs_3-5 224 8 9.18
glass-0-4_vs_5 92 10 9.22
ecoli-0-3-4-6_vs_5 205 8 9.25
ecoli-0-3-4-7_vs_5-6 257 8 9.28
vowel 988 14 9.98
ecoli-0-6-7_vs_5 220 7 10
glass-0-1-6_vs_2 192 10 10.29
ecoli-0-1-4-7_vs_2-3-5-6 336 8 10.59
glass-0-6_vs_5 108 10 11
glass-0-1-4-6_vs_2 205 10 11.06
glass2 214 10 11.59
ecoli-0-1-4-7_vs_5-6 332 7 12.28
cleveland-0_vs_4 177 14 12.62
shuttle-c0-vs-c4 1829 10 13.87
yeast-1_vs_7 459 8 14.3
glass4 214 10 15.46
ecoli4 336 8 15.8
page-blocks-1-3_vs_4 472 11 15.86
abalone 731 9 16.4
glass-0-1-6_vs_5 184 10 19.44
yeast-1-4-5-8_vs_7 693 9 22.1
yeast4 1484 9 28.1
yeast128 947 9 30.57
yeast5 1484 9 32.73
winequality-red-8_vs_6 656 12 35.44
ecoli_013vs26 281 8 39.14
abalone-17_vs_7-8-9-10 2338 9 39.31
yeast6 1483 9 41.37
winequality-white-3_vs_7 900 12 44
winequality-red-8_vs_6-7 855 12 46.5
kddcup-land_vs_portsweep 1060 39 49.48
abalone-19_vs_10-11-12-13 1622 9 49.69
winequality_white 1481 12 58.24
poker-8-9_vs_6 1484 11 58.36
winequality-red-3_vs_5 691 12 68.1
abalone_20 1916 8 72.69
kddcup-land_vs_satan 1609 39 79.45
poker-8-9_vs_5 2074 11 81.96
poker_86 1477 11 85.88
kddr_rookkit 2225 42 100.14
Table 3: Summary of the datasets for multiclass classification tasks
Dataset # Features # Classes No of Samples/Class
thyroid 21 3 $6$, 368, 166
new-thyroid 5 3 150, 35, 30
contraceptive 9 3 629, 511, 333
shuttle 9 3 1706, 338, 131
Dry_Bean_Dataset 16 7 3546, 2636, 2027, 1928, 1630, 1322, 522
balance 4 3 288, 288, 49
pageblocks 10 5 4913, 329, 115, 88, 28
HCV 12 4 540, 30, 24, 21
yeast 8 10 463, 429, 244, 163, 51, 44, 35, 30, 20, 5
wine 13 3 71, 59, 48
4.2 Experimental Setup
To ensure proper validation and avoid data leakage, the data was first split into training and testing folds. The algorithms are applied only to the training set and the performance is evaluated on the testing set. A repeated stratified K-fold cross-validation strategy (5 folds, 10 repeats) was adopted to ensure better generalization. The average of the results from all different testing folds (50) is considered as the performance measure.
Before training, the data was normalized using z-score (StandardScaler from sklearn library). We experimented with 4 different classification algorithms in this study: LR, SVM, DT, and Random Forest (RF). The default parameters of the scikit-learn library were utilized to implement these classifiers. None of the parameters (except ’class_weight) of the base classifiers were tuned to ensure a valid comparison.
In the case of the proposed iCost algorithm, only the ’cost-factor’ parameter was tuned using the grid-search technique. We used type = ’ins’ to obtain results for the proposed modified CS approach. The parameter setting for the grid-search implementation is provided in Table 4. The MCC score was utilized as the scoring criteria. The performance from other different types of categorization available in the framework may vary slightly. However, those are not included in this manuscript.
Table 4: Parameter settings for the grid-search implementation
Parameter Value Cost-factor parameter setting
Algorithm original 0.8*IR, 0.9*IR, IR, 1.1*IR, 1.2*IR
Algorithm Neighborhood ’pure’ : [1, 0.2*IR]
’safe’: [0.25*IR, 0.35*IR, 0.5*IR]
’border’: [0.75*IR, 0.9*IR, IR, 1.1*IR, 1.25*IR]
Algorithm MST ’linked’ : [0.75*IR, 0.9*IR, IR, 1.1*IR, 1.25*IR]
’normal’: [0.3*IR, 0.5*IR, 0.7*IR]
4.3 Performance Metrics
Assessing the performance of different techniques on skewed data can be challenging [1]. Eight different performance metrics were computed for evaluation in this study: MCC, ROC-AUC, G-mean, F1-score, sensitivity, specificity, precision, and accuracy. ML algorithms often excel at predicting instances from the majority class but tend to perform poorly on the minority class. Consequently, traditional performance metrics like accuracy can be misleading because they do not account for the distribution of classes.
To give an example, if 95% of the data belongs to one class and 5% to another, a model that always predicts the majority class will have 95% accuracy but will be useless for predicting the minority class. Sensitivity and specificity are two class-specific metrics that manifest the performance accuracy of the minority and majority classes, respectively. They only show the performance of a particular class. As a result, it is difficult to apprehend the entire performance spectrum from these measures.
Composite metrics are more suitable for performance measurements on imbalanced data [47, 48]. The g-mean score exhibits a broader picture by providing the geometric mean of sensitivity and specificity. F1-score provides the harmonic mean between sensitivity and precision. The algorithm needs to attain a high sensitivity to have a better G-mean or F1-score. While these two metrics are quite popular and have their advantages, there are certain issues and limitations associated with them. They do not consider the actual number of misclassifications.
To elaborate, a sensitivity score of 0.8, observed with only 100 minority-class samples in the dataset, implies that the model misclassified 20 instances from the minority class. On the other hand, a specificity score of 0.8 with 10,000 majority-class samples in the same data indicates that the model misclassified 2,000 instances. This is a huge difference in the number of misclassifications which is not apparent from class-specific metrics. An algorithm may improve the sensitivity score to 0.9 which indicates that only 10 minority class instances have been misclassified. However, improving sensitivity often comes at the cost of lowering the specificity score. If the specificity score is reduced to 0.7 by the algorithm, then the model is now making 3000 misclassifications on the majority class. For both cases, the overall g-mean score remains almost the same. However, the model is now misclassifying around 1000 more samples which is not captured by these metrics.
To be more vigilant regarding performance measurements, we consider other metrics such as MCC and ROC-AUC [49]. MCC provides a robust performance measure as it considers the actual number of misclassifications for both classes [50]. This comprehensive consideration provides a balanced measure of model performance. The MCC score is high only when the classifier performs well across all cases, making it one of the most robust measures of classification performance [51]. However, as stated in several previous literature [47, 52], the performance of algorithms on imbalanced data cannot be sufficiently represented by a single metric. Therefore, we take into consideration four composite metrics to understand how the proposed approach fares compared to other techniques used in imbalanced learning.
For multiclass settings, three different averaging criteria are available: macro, micro, and weighted [53]. The macro average is usually considered the best for imbalanced datasets as it will give a lower score if the model performs poorly on less frequent classes, even if the majority class is well-predicted. Therefore, the macro-average results are calculated and presented in this manuscript.
4.4 Performance Comparison
One of the main objectives of the developed methodology is to achieve better generalization and performance over the traditional CS implementations of different ML algorithms. As such, we compared the performance of the proposed approach with the traditional CS learning techniques to evaluate the differences.
There are many other techniques available that are used in imbalanced learning. Data resampling is a completely different paradigm in imbalanced data handling. As such, comparing performance with state-of-the-art sampling techniques is irrelevant since that is not the primary objective of the proposed algorithm. Nevertheless, we compare our results with some of the popular sampling techniques. These include oversampling approaches such as SMOTE [31], ADASYN [54], Borderline-SMOTE (BLS) [55], and Random oversampling (ROS); undersampling techniques such as Random undersampling (RUS), Edited nearest neighbor (ENN), and Neighborhood cleaning (NC); hybrid sampling technique SMOTE-Tomek. Details of the performance measures from these approaches have been reported in the following section. The imblearn library was utilized to implement these sampling techniques using the default parameter settings [56].
5 Results and Discussion
In this section, we present and discuss the results we have obtained during the experiment. We have measured the performance of 4 different classifiers on 65 binary imbalanced datasets and 10 multiclass imbalanced datasets using 8 different metrics. All these measures cannot be included here due to space constraints. They are provided in separate supplementary files. We mainly considered the MCC, ROC-AUC, G-mean, and F1-score for comparison. The results from the ’neighborhood’ categorization and type = ’ins’ are elaborately presented here and have also been applied to the multiclass scenario.
5.1 Performance from traditional approaches
Standard classification algorithms do not perform well on imbalanced datasets. The performance gets even worse with higher imbalances and class overlapping [30]. In many cases, the g-mean score was found to be 0 (sensitivity=0, specificity=100), indicating a clear bias towards the majority class. CSL provides a good improvement in performance as can be observed from Fig 6. It shows the average G-mean scores obtained on 65 datasets. Among the classifiers, CS-SVM provided the highest G-mean score. Improvement in performance was the maximum in the LR classifier. DT is found to be less sensitive to CS approaches. For the RF classifier, the average performance actually dropped a little when CSL was applied.
One thing to note here is that this is the average g-mean score obtained on the 65 datasets we tested. It does not portray the individual cases. In many of them, the performance improved from the CS-RF or CS-DT approaches. While in many others the performance dropped and on average, there was a decline in the overall g-mean score. The average results from other metrics are provided in Table 5, Table 6, Table 7, and Table 8.
Figure 6: Performance comparison among standard and cost-sensitive approaches.
5.2 Performance comparison of the proposed iCost algorithm with the standard CS approach
In this study, we propose a modification to the original implementation of CSL by introducing an instance complexity-based CS framework. In our approach, different misclassification costs are assigned to minority-class instances depending on their difficulty level. The samples located near the border are highly penalized compared to the samples that are away from the decision boundary or surrounded by instances of the same class. This ensures that the safe samples do not overshadow other majority-class instances and create unwarranted bias. Thus, our proposed approach provides a more plausible CSL framework where instances are weighted according to their complexity, rather than indiscriminately. This modification provides quite an improvement in performance.
Figure 7: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the LR classifier on 65 datasets.
Figure 8: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the SVM classifier on 65 datasets.
Figure 9: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the DT classifier on 65 datasets.
Figure 10: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the RF classifier on 65 datasets.
The changes in the MCC score obtained from the proposed algorithm as compared to the traditional CS approach are presented in Fig. 7, Fig. 8, Fig. 9, and Fig. 10 for the LR, SVM, DT, and RF classifiers, respectively. The datasets are sorted based on IR in ascending order on the x-axis. Other measures of performance on individual datasets are provided as supplementary files in the GitHub repository.
As can be observed from Fig. 7, a significant improvement is noticeable in most datasets for the LR classifier. For a few datasets, there was no change in performance and for only one dataset, the performance dropped slightly. As for the SVM classifier, performance also improved in most datasets. However, the increment was lower than the LR classifier. In several of the datasets, the performance remained unchanged. This happened mostly in highly imbalanced cases. In such cases, the number of minority class samples is usually limited and they are surrounded by instances of the opposite class, leaving very few pure and safe samples. As a result, almost all the samples are border samples which are weighted equally and the algorithm works the same as the standard CS approach. In the case of the DT classifier, there was a deterioration in performance in some of the datasets. However, in most cases, it was almost negligible. As for the other datasets, there was a decent improvement in performance. For the RF classifier, there was an overall improvement in most cases. Performance dropped marginally for a couple of datasets.
Overall, the most improvement was noticeable in the LR classifier (3.3% on average per dataset). For the other three classifiers, the improvement was around 2% on average. MCC is a highly reliable performance metric, and an improvement in MCC indicates that our proposed method effectively reduces the number of misclassifications. Fig. 11 demonstrates the other performance measures (average) obtained on 65 imbalanced datasets for the LR classifier. As can be observed from the figure, there is an increase in performance from the proposed algorithm over the traditional CS approach for almost all the measures. Only the sensitivity score dropped marginally. The number of minority class samples available in the datasets is usually very small. As a result, only a few misclassifications lead to a large change in the sensitivity score. However, in terms of all four composite metrics, the proposed algorithm achieved better results compared to the traditional CS approach. For the RF classifier, both sensitivity and specificity scores improved by employing the proposed framework. Similar improvements are noticeable for other classifiers.
Figure 11: Average performance measures on 65 datasets for the LR classifier.
5.3 Performance comparison of the proposed algorithm with other sampling techniques
The performance was also compared with some of the popular sampling approaches. Data resampling techniques undertake a completely different path in addressing the class imbalance and they have been widely applied in imbalanced learning tasks. The average of performance measures obtained from these approaches on the 65 imbalanced datasets are provided in Table 5, Table 6, Table 7, and Table 8 for the LR, SVM, DT, and RF classifiers, respectively.
As can be observed from the tables, the proposed framework usually produced better prediction performance compared to the other approaches in terms of precision, ROC-AUC, G-mean, MCC, and F1-score. When the data was resampled using SMOTE, the most popular sampling technique, G-mean scores of 77.62% and 78.45% were obtained for the LR and SVM classifiers, respectively. Our proposed approach outperformed this well-established method by producing G-mean scores of 81.03% and 82.8%, respectively. A good improvement in performance was noticeable compared to other approaches as well.
For the RF classifier, however, the sampling techniques performed quite well compared to the CSL approach. While the iCost algorithm outperformed the traditional CS-RF approach, it could not outdo other sampling approaches. RF is an ensemble method that works by creating bootstrap subsets. In each bootstrap subset, the IR is different than the original dataset. The CS algorithms do not capture that as it uses the overall IR. That is why the performance gets constantly lower compared to the sampling techniques. We plan to address this issue in the future.
For all cases, the standard classifiers provided the highest specificity score due to the bias towards the majority class. For the SVM classifier, RUS provided the highest sensitivity score. However, RUS also attained the lowest specificity score. In RUS, a large number of majority-class samples are removed from the data to attain balance. Consequently, this approach provided the best performance for the minority class at the cost of the lowest performance for the majority class. This makes the prediction framework quite unreliable. A higher sensitivity score does not really indicate that the approach is actually performing well. Metrics such as ROC-AUC or MCC score manifests a more clear picture and the RUS approach produced a MCC score of 48.15%. Compared to that, the iCost algorithm achieved a much higher MCC score of 60.02%.
Table 5: Performance measures obtained from different approaches for the LR classifier
Metrics LR SMOTE ADASYN BLS ROS RUS ENN NC S_Tomek CS-LR iCost (Proposed)
Sensitivity 18.34 79.88 80.81 78.37 80.02 80.19 25.83 25.91 79.98 80.21 78.43
Specificity 99.16 80.46 78.16 80.77 79.70 75.16 96.40 96.30 80.44 83.26 85.47
Precision 33.64 46.70 42.75 45.56 45.46 41.65 36.05 37.31 46.69 41.87 47.13
ROC-AUC 58.75 80.17 79.48 79.57 79.86 77.67 61.11 61.10 80.21 81.73 81.95
G-mean 24.79 77.62 76.57 76.11 76.87 74.48 30.79 30.95 77.63 80.29 81.03
MCC 21.72 49.50 46.80 48.72 48.58 44.18 25.96 26.21 49.53 48.2 51.56
F1-score 21.81 52.47 49.94 51.90 51.53 47.50 27.83 28.10 52.46 50.57 53.96
Table 6: Performance measures obtained from different approaches for the SVM classifier
Metrics SVM SMOTE ADASYN BLS ROS RUS ENN NC S_Tomek CS-SVM iCost(Proposed)
Sensitivity 41.21 77.28 78.64 74.10 77.19 83.37 51.32 51.64 77.30 78.12 76.53
Specificity 97.80 86.50 84.33 87.31 85.19 76.62 94.26 94.54 86.50 89.2 91.41
Precision 53.39 55.64 52.47 56.43 54.41 44.96 53.84 55.45 55.64 55.27 58.32
G-mean 48.47 78.45 78.03 76.05 77.70 76.43 54.78 55.25 78.46 81.05 82.8
MCC 42.99 56.13 54.05 55.59 54.80 48.15 47.24 48.23 56.14 57.99 60.02
ROC-AUC 69.51 81.89 81.49 80.70 81.19 79.99 72.79 73.09 81.90 83.66 83.97
F1-Score 43.98 58.07 56.26 57.95 57.03 50.57 49.03 49.86 58.07 60.03 61.97
Table 7: Performance measures obtained from different approaches for the DT classifier
Metrics DT SMOTE ADASYN BLS ROS RUS ENN NC S_Tomek CS-DT iCost(Proposed)
Sensitivity 55.46 62.18 63.38 60.07 53.73 81.81 61.39 62.48 62.20 54.69 56.1
Specificity 93.61 92.02 91.69 92.59 94.50 74.48 90.44 90.62 92.00 96 96.1
Precision 53.60 51.95 52.15 53.05 55.00 36.37 50.31 51.54 51.78 57.38 58.36
G-mean 65.34 70.53 70.80 68.46 63.22 76.29 68.58 68.71 70.47 65.14 69.91
MCC 48.32 50.04 50.49 49.83 48.29 41.07 48.24 49.57 49.92 51.39 52.73
ROC-AUC 74.54 77.10 77.53 76.33 74.12 78.14 75.92 76.55 77.10 75.05 76.1
F1-Score 51.87 53.45 53.71 53.21 51.19 44.09 52.16 53.12 53.34 54.68 55.9
Table 8: Performance measures obtained from different approaches for the RF classifier
Metrics RF SMOTE ADASYN BLS ROS RUS ENN NC S_Tomek CS-RF iCost(Proposed)
Sensitivity 51.10 64.42 64.53 61.20 56.60 84.38 58.18 58.35 64.52 49.22 50.38
Specificity 98.51 96.38 95.94 96.62 97.69 82.78 96.12 96.26 96.35 98.63 98.66
Precision 66.99 64.07 62.41 65.35 67.47 41.70 63.76 64.43 64.08 66.27 67.35
G-mean 63.56 72.93 72.82 69.95 65.36 82.73 64.64 65.01 72.99 61.63 62.60
MCC 54.99 60.28 59.26 59.18 58.04 49.46 56.62 57.10 60.32 53.60 54.76
ROC-AUC 74.81 80.40 80.24 78.91 77.15 83.58 77.15 77.31 80.43 73.92 74.52
F1-Score 55.61 62.80 61.91 61.48 59.43 50.63 58.42 58.82 62.84 54.08 55.18
5.4 Performance in multiclass datasets
The proposed algorithm was also tested in multiclass imbalanced scenarios. A total of 10 datasets with different numbers of classes and different categories (multi-minority, multi-majority, or equal ratio) were utilized in the experiment. We compared the performance of the iCost framework with the traditional CS approach under the OvR setting. Detailed performance measures on individual datasets are provided in the supplementary files. The average performance measures are reported in Table 9. The changes in the MCC score from the proposed algorithm compared to standard CS techniques are illustrated in Fig. 12, Fig. 13, Fig. 14, and Fig. 15 for the LR, SVM, DT, and RF classifiers, respectively.
Multiclass classification is naturally more complicated compared to binary classifications [57]. There are many classes involved and correctly identifying the instances in the minority classes becomes more difficult. Prior to discussing the results, one thing to consider in multiclass scenarios is that even a small change in the MCC or G-mean score is hard to achieve given that there are many classes and the MCC score takes into account predictions in all categories. Achieving higher scores indicates improving correct predictions in minority classes without causing misclassifications of other class instances.
As can be observed from Table 9, the proposed algorithm provides quite a good improvement in overall performance over the standard CS approach. Both the sensitivity and specificity scores improved, indicating the effectiveness of the iCost framework. The precision score along with the F1 and G-mean scores also improved. For the LR classifier, there was almost a 4% improvement in the sensitivity, MCC, and F1 score. Similar improvement was also observed for the SVM classifier. For the DT classifier, the average improvement was very small. However, for none of these cases, there was any decline in the overall scores. Only for the RF classifier, there is a marginal decline in performance. But it is less than 0.05%, making it inconsequential. The drop in performance was mostly from only one dataset, while in other cases, the performance remained almost unchanged.
Table 9: Average performance measures on 10 multiclass datasets
Metrics Accuracy Sensitivity Specificity Precision G-mean MCC F1-Score
LR 82.51 57.86 85.39 69.74 69.28 56.21 59.69
CS-LR 83.22 66.75 88.16 75.20 76.33 61.95 67.88
iCost-LR 84.61 70.27 88.93 77.92 78.56 64.95 71.07
SVM 86.24 68.58 89.44 75.89 77.78 67.31 70.00
CS-SVM 85.70 76.22 92.45 78.28 83.57 70.68 75.63
iCost-SVM 87.39 77.91 93.49 79.87 85.00 73.46 77.76
DT 82.87 75.31 93.44 75.46 83.15 71.25 74.40
CS-DT 82.82 75.37 93.50 75.32 83.24 71.15 74.23
iCost-DT 83.07 75.72 93.65 75.64 83.54 71.73 74.58
RF 87.86 79.58 94.48 81.59 86.26 78.45 80.06
CS-RF 87.98 79.74 94.56 81.48 86.39 78.66 80.03
iCost-RF 87.95 79.67 94.55 81.43 86.35 78.62 79.96
Figure 12: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the LR classifier on 10 multiclass datasets.
Figure 13: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the SVM classifier on 10 multiclass datasets.
Figure 14: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the DT classifier on 10 multiclass datasets.
Figure 15: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the RF classifier on 10 multiclass datasets.
To observe the performance on individual datasets, we can look at Fig. 12, Fig. 13, Fig. 14, and Fig. 15. For the LR classifier, there is a good improvement in performance in most datasets. In the new thyroid dataset, the MCC score improved by around 12%. For the SVM classifier, performance improved significantly in 3 datasets. For others, the improvement was around 1%. For the DT classifier, the performance dropped slightly (0.01%) in 2 datasets. In other cases, the performance improved. For the RF classifier, in the ’balance’ dataset, the MCC score dropped by 1.5%. In most other cases, the changes in the MCC score were very limited. Given the nature of the ensemble algorithms, they provide well-generalized performance. That is why the changes are limited in such an ensemble algorithm.
Overall, our proposed modification in the CSL framework works equally in multiclass classification scenarios as well. Good improvement in performance can be observed, especially in classifiers such as SVM and LR. Ensemble algorithms (such as RF classifiers), which are more robust, are less susceptible to such modifications.
5.5 Performance from MST categorization criteria
Besides the ’neighborhood’ approach, the ’MST’ categorization criterion was also tested on the 65 imbalanced datasets. The detailed performance measures are provided as supplementary files in the GitHub repository. There is an overall improvement in performance over the traditional CS framework from this categorization as well. The changes in the MCC score from the LR classifier are illustrated in Fig. 16. Similar improvements in performance as the ’neighborhood’ approach are noticeable for other classifiers as well.
It is difficult to say which categorization criteria between the two works better. In some datasets, the neighborhood categorization works better, while in others, MST works better. For the LR classifier, in terms of MCC score, the neighborhood categorization gives better performance than the MST categorization in 29 datasets out of 65. For the SVM classifier, however, in only 13 datasets, neighborhood categorization works better. For the DT and RF classifiers, the numbers are 38 and 50, respectively. The difference in performance in many cases is quite negligible.
The difference in the average MCC scores from these approaches on 65 imbalanced datasets is illustrated in Fig. 17. As can be observed from the figure, there is an overall improvement in performance from the proposed approach over the standard CS implementations of the 4 different classification algorithms. There is a slight difference between the two categorization criteria.
Figure 16: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the LR classifier on 65 imbalanced datasets for MST categorization.
Figure 17: Difference in the average MCC scores between the traditional CS method and the proposed iCost approaches
6 Conclusion
In this study, a modified cost-sensitive learning framework is proposed where data difficulty factors are taken into consideration while penalizing the instances. In the traditional framework, all the instances are weighted equally. This is impractical and can bias the prediction toward the minority-class, leading to an increased amount of false positives. Uniform instance weighting overfits the data by unusually deforming the decision boundary. This does not fare well during testing, resulting in a higher number of misclassifications. To alleviate the issue, we weigh the minority-class samples in the overlapping region more strongly compared to the ones in the non-overlapping region. Carefully assigning higher weights to the hard-to-learn examples while reducing the weights of the other provides a more plausible weighting mechanism, resulting in fewer misclassifications.
We proposed two different instance categorization criteria in this manuscript. We have tested our algorithm on 65 binary and 10 multiclass imbalanced datasets using 4 different classifiers and observed an increment in performance in most cases. The proposed algorithm is computationally light similar to traditional CS approaches. The modification improves the original CS framework implemented in popular Python libraries such as scikit-learn by introducing some reasonable adjustments to it.
Further improvement is possible in the proposed algorithm and we plan to explore those in future works. We measured the instance complexity based on nearest neighbors or MST. However, there are other data complexity measures such as local sets [17] which can be considered. We plan to evaluate the possibility of incorporating those into the framework. While we worked with four different classifiers in this study, the proposed algorithm can be used with other classification algorithms such as XGBoost. The default values for the proposed approach are empirically set. Further research is required in this regard to understand how different data difficulty factors are related to the cost-factor values.
To conclude, this research introduces a novel approach in the cost-sensitive learning domain for imbalanced classification tasks. Penalizing minority-class instances according to their complexity provides improved prediction performance in both binary and multiclass classification scenarios. The implementation details as well as the code are provided to ensure reproducability. The code is written in Python ensuring full compatability with the scikit-learn library. The work presents a new perspective in the field of imbalanced domains and paves the way for the development of new strategies that are based on data-specific characteristics.
7 Appendix
Supplementary files and codes are available in this repository:
https://github.com/newaz-aa/iCost